Solve and Check the following rational equations

4 - 1 = 1
- - -
x+2 x x

Answer by stanbon(48558)

Solve and Check the following rational equations

4 - 1 = 1
- - -
x+2 x x
======================
Multiply thru by x(x+2) to get:
4x - (x+2) = (x+2)
4x = 2(x+2)
4x = 2x + 4
2x = 4
x = 2
===============
Ceers,
Stan H.